# Improved SINR Approximation for Downlink SDMA-based Networks with Outdated Channel State Information

### v1

@@ V Conclusion
Using an RSMA-based MU-MIMO downlink scenario for benchmarking, we demonstrated that the proposed approximation is particularly effective in massive MIMO regimes (i.e., $N_{\rm t}\gg K$), providing significantly tighter estimates of the ergodic sum-rate and mitigating the optimistic bias introduced by existing models.
@@ Appendix A Proof of Lemma 1
$\mathbb{E}[X]=K(1-\epsilon^{2})+(N_{\rm t}-K+1)\epsilon^{2}.$ (20)
@@
$\zeta_{1}$ $=\mathbb{E}\left\{\left((1-\epsilon^{2})\sum_{j\in\mathcal{K}}|\boldsymbol{e}_{k}^{\rm{H}}[m]\boldsymbol{p}_{j}|^{2}\right)^{2}\right\},$ (22)
$\zeta_{2}$ $=\mathbb{E}\left\{\left(\epsilon^{2}|\boldsymbol{h}_{k}^{\rm{H}}[m-1]\boldsymbol{p}_{k}|^{2}\right)^{2}\right\},$ (23)
$\zeta_{3}$ $=\mathbb{E}\left\{2(1-\epsilon^{2})\sum_{j\in\mathcal{K}}|\boldsymbol{e}_{k}^{\rm{H}}[m]\boldsymbol{p}_{j}|^{2}\epsilon^{2}|\boldsymbol{h}_{k}^{\rm{H}}[m-1]\boldsymbol{p}_{k}|^{2}\right\},$ (24)
$\zeta_{4}$ $=\mathbb{E}\left\{(2\sqrt{\epsilon^{2}(1-\epsilon^{2})}\Re\{\boldsymbol{h}_{k}^{\rm{H}}[m-1]\boldsymbol{p}_{k}\boldsymbol{e}_{k}^{\rm{H}}\boldsymbol{p}_{k})\}^{2}\right\}.$ (25)
@@
$\zeta_{1}$ $=K(1-\epsilon^{2})^{2}+K^{2}(1-\epsilon^{2})^{2},$ (26)
$\zeta_{2}$ $=(N_{\rm t}-K+1)\epsilon^{4}+(N_{\rm t}-K+1)^{2}\epsilon^{4}.$ (27)
To derive expressions for $\zeta_{3}$ and $\zeta_{4}$, we first perform some algebraic manipulations to obtain (28) and (29), respectively, at the top of next page. In the following derivations, the components of the vectors $\boldsymbol{h}_{k}$, $\boldsymbol{p}_{k}$, and $\boldsymbol{e}_{k}$ are denoted as scalars with an additional subindex. Also, for the sake of compactness, the following ancillary definitions are made: $\phi[m]=\mathbb{E}\left\{\left(\sum_{j=1}^{N_{\rm t}}h_{k,j}[m-1]\overline{p}_{k,j}\right)^{2}\right\}$ and $\xi[m]=\sum_{j=1}^{N_{\rm t}}\mathbb{E}\left\{|h_{k,j}[m-1]|^{2}|p_{k,j}|^{2}\right\}$.
$\zeta_{3}=2(1-\epsilon^{2})\epsilon^{2}\mathbb{E}\left\{\sum_{j=1}^{K}\sum_{a=1}^{N_{\rm t}}\sum_{b=1}^{N_{\rm t}}\sum_{c=1}^{N_{\rm t}}\sum_{d=1}^{N_{\rm t}}\overline{e}_{k,a}[m]p_{j,a}e_{k,b}[m]\overline{p}_{j,b}\overline{h}_{k,c}[m-1]p_{k,c}h_{k,d}[m-1]\overline{p}_{k,d}\right\},$ (28)
$\zeta_{4}=2(1-\epsilon^{2})\epsilon^{2}\mathbb{E}\left\{\sum_{a=1}^{N_{\rm t}}\sum_{b=1}^{N_{\rm t}}\sum_{c=1}^{N_{\rm t}}\sum_{d=1}^{N_{\rm t}}\overline{h}_{k,a}[m-1]p_{k,a}\overline{e}_{k,b}[m]p_{k,b}h_{k,c}[m-1]\overline{p}_{k,c}e_{k,d}[m]\overline{p}_{k,d}\right\},$ (29)
@@
$c_{1}=2(1-\epsilon^{2})\epsilon^{2}\sum_{j=1}^{K}\sum_{a=1}^{N_{\rm t}}\mathbb{E}\{|e_{k,a}[m]|^{2}|p_{j,a}|^{2}\\ |h_{k,a}[m-1]|^{2}|p_{k,a}|^{2}\}$ (30)
Assuming independence between the error and precoding components, and also between the outdated channel and the precoding vector components for $j\neq k$, we can reexpress:
$c_{1}\approx 2(1-\epsilon^{2})\epsilon^{2}\sum_{j=1}^{K}\sum_{a=1}^{N_{\rm t}}\mathbb{E}\left\{|e_{k,a}[m]|^{2}\right\}\mathbb{E}\left\{|p_{j,a}|^{2}\right\}\\ \times\mathbb{E}\left\{|h_{k,a}[m-1]|^{2}|p_{k,a}|^{2}\right\}.$ (31)
Noting that $\mathbb{E}\left\{|e_{j,l}[m]|^{2}\right\}=1$ and $\mathbb{E}\left\{|p_{j,l}|^{2}\right\}\approx 1/N_{\rm t}$, $\forall j,l$, through mathematical manipulation, we have:
$c_{1}\approx 2(1-\epsilon^{2})\epsilon^{2}K\phi[m]/N_{\rm t}.$ (32)
Following a similar rationale, the cases for $c_{2}$ to $c_{5}$ can be computed as summarized in Table I.
@@
$a=b=c$ $c_{2}\approx 2(1-\epsilon^{2})\epsilon^{2}K\left(\phi[m]-\xi[m]\right)/N_{\rm t}.$
$a=b=d$ $c_{3}\approx 2(1-\epsilon^{2})\epsilon^{2}K\left(\phi[m]-\xi[m]\right)/N_{\rm t}.$
$a=b$ and $c=d$ $c_{4}\approx 2(1-\epsilon^{2})\epsilon^{2}K(N_{\rm t}-1)\xi[m]/N_{\rm t}.$
$a=b$ $c_{5}\approx 2(1-\epsilon^{2})\epsilon^{2}K(N_{\rm t}-2)(\phi[m]-\xi[m])/N_{\rm t}.$
@@
$\zeta_{3}\approx 2(1-\epsilon^{2})\epsilon^{2}K\phi[m].$ (33)
@@
$\zeta_{4}\approx 2(1-\epsilon^{2})\epsilon^{2}\phi[m].$ (34)
@@
$\mathrm{Var}\left\{X\right\}=\epsilon^{4}(N_{\rm t}+1)+(1-2\epsilon^{2})K+\mu,$ (36)
where $\mu$ is:
$\mu=2\epsilon^{2}(1-\epsilon^{2})(N_{\rm t}-K+1).$ (37)
This completes the proof.

### v2

@@ V Conclusion
Using an RSMA-based MU-MIMO downlink scenario for benchmarking, we demonstrated that the proposed approximation is particularly effective in massive MIMO regimes (i.e., $N_{\rm t}\gg K$), providing significantly tighter estimates of the ergodic sum-rate and mitigating the optimistic bias introduced by existing models. Future work may extend the proposed analysis to non-Rayleigh fading environments by incorporating different propagation conditions, or to consider alternative precoding schemes such as regularized ZF or MMSE.
@@ Appendix A Proof of Lemma 1
$\mathbb{E}[X]=(N_{\rm t}-K+1)\epsilon^{2}\alpha_{k}+(1-\alpha_{c})(1-\epsilon^{2}).$ (20)
@@
$\zeta_{1}$ $=\mathbb{E}\left\{\left((1-\epsilon^{2})\sum_{j\in\mathcal{K}}\alpha_{j}|\mathbf{e}_{k}^{\rm{H}}[m]\mathbf{p}_{j}|^{2}\right)^{2}\right\},$ (22)
$\zeta_{2}$ $=\mathbb{E}\left\{\left(\epsilon^{2}\alpha_{k}|\mathbf{h}_{k}^{\rm{H}}[m-1]\mathbf{p}_{k}|^{2}\right)^{2}\right\},$ (23)
$\zeta_{3}$ $=\mathbb{E}\left\{2(1-\epsilon^{2})\epsilon^{2}\alpha_{k}\sum_{j\in\mathcal{K}}\alpha_{j}|\mathbf{e}_{k}^{\rm{H}}[m]\mathbf{p}_{j}|^{2}|\mathbf{h}_{k}^{\rm{H}}[m-1]\mathbf{p}_{k}|^{2}\right\},$ (24)
$\zeta_{4}$ $=\mathbb{E}\left\{\left(2\sqrt{\epsilon^{2}(1-\epsilon^{2})}\alpha_{k}\Re\{\mathbf{h}_{k}^{\rm{H}}[m-1]\mathbf{p}_{k}\mathbf{e}_{k}^{\rm{H}}\mathbf{p}_{k}\}\right)^{2}\right\}.$ (25)
@@
$\zeta_{1}$ $=\sum_{j\in\mathcal{K}}\alpha_{j}^{2}(1-\epsilon^{2})^{2}+(1-\epsilon^{2})^{2}(1-\alpha_{c})^{2},$ (26)
$\zeta_{2}$ $=(N_{\rm t}-K+1)\epsilon^{4}\alpha_{k}^{2}+(N_{\rm t}-K+1)^{2}\epsilon^{4}\alpha_{k}^{2}.$ (27)
To derive expressions for $\zeta_{3}$ and $\zeta_{4}$, we first perform some algebraic manipulations to obtain (28) and (29), respectively, at the top of next page. In the following derivations, the components of the vectors $\mathbf{h}_{k}$, $\mathbf{p}_{k}$, and $\mathbf{e}_{k}$ are denoted as scalars with an additional subindex. For the sake of compactness, the following ancillary definitions are used in the derivations: $\phi[m]=\mathbb{E}\left\{\left(\sum_{j=1}^{N_{\rm t}}h_{k,j}[m-1]\overline{p}_{k,j}\right)^{2}\right\}$, and $\xi[m]=\sum_{j=1}^{N_{\rm t}}\mathbb{E}\left\{|h_{k,j}[m-1]|^{2}|p_{k,j}|^{2}\right\}$.
$\zeta_{3}=2(1-\epsilon^{2})\epsilon^{2}\alpha_{k}\mathbb{E}\left\{\sum_{j=1}^{K}\sum_{a=1}^{N_{\rm t}}\sum_{b=1}^{N_{\rm t}}\sum_{c=1}^{N_{\rm t}}\sum_{d=1}^{N_{\rm t}}\alpha_{j}\overline{e}_{k,a}[m]p_{j,a}e_{k,b}[m]\overline{p}_{j,b}\overline{h}_{k,c}[m-1]p_{k,c}h_{k,d}[m-1]\overline{p}_{k,d}\right\},$ (28)
$\zeta_{4}=2(1-\epsilon^{2})\epsilon^{2}\alpha_{k}^{2}\mathbb{E}\left\{\sum_{a=1}^{N_{\rm t}}\sum_{b=1}^{N_{\rm t}}\sum_{c=1}^{N_{\rm t}}\sum_{d=1}^{N_{\rm t}}\overline{h}_{k,a}[m-1]p_{k,a}\overline{e}_{k,b}[m]p_{k,b}h_{k,c}[m-1]\overline{p}_{k,c}e_{k,d}[m]\overline{p}_{k,d}\right\},$ (29)
@@
$c_{1}=2(1-\epsilon^{2})\epsilon^{2}\alpha_{k}\sum_{j=1}^{K}\sum_{a=1}^{N_{\rm t}}\mathbb{E}\{\alpha_{j}|e_{k,a}[m]|^{2}|p_{j,a}|^{2}\\ \times|h_{k,a}[m-1]|^{2}|p_{k,a}|^{2}\}.$ (30)
Assuming independence between the error and precoding components, and also between the outdated channel and the precoding vector components for $j\neq k$, and because the power allocation coefficient for each user is constant, we can reexpress:
$c_{1}\approx 2(1-\epsilon^{2})\epsilon^{2}\alpha_{k}\sum_{j=1}^{K}\sum_{a=1}^{N_{\rm t}}\alpha_{j}\mathbb{E}\left\{|e_{k,a}[m]|^{2}\right\}\mathbb{E}\left\{|p_{j,a}|^{2}\right\}\\ \times\mathbb{E}\left\{|h_{k,a}[m-1]|^{2}|p_{k,a}|^{2}\right\}.$ (31)
Noting that $\mathbb{E}\left\{|e_{j,l}[m]|^{2}\right\}=1$, $\mathbb{E}\left\{|p_{j,l}|^{2}\right\}\approx 1/N_{\rm t}$, $\forall j,l$, and $\sum_{j\in\mathcal{K}}\alpha_{j}=(1-\alpha_{c})$ through mathematical manipulation, we have:
$c_{1}\approx 2(1-\epsilon^{2})\epsilon^{2}\alpha_{k}(1-\alpha_{c})\phi[m]/N_{\rm t}.$ (32)
Following a similar rationale, the cases for $c_{2}$ to $c_{5}$ can be computed as summarized in Table I, where $\kappa=2(1-\epsilon^{2})\epsilon^{2}\alpha_{k}(1-\alpha_{c})$}.
@@
$c_{2}$ $a=b=c$ $\kappa\left(\phi[m]-\xi[m]\right)/N_{\rm t}.$
$c_{3}$ $a=b=d$ $\kappa\left(\phi[m]-\xi[m]\right)/N_{\rm t}.$
$c_{4}$ $a=b$ and $c=d$ $\kappa(N_{\rm t}-1)\xi[m]/N_{\rm t}.$
$c_{5}$ $a=b$ $\kappa(N_{\rm t}-2)(\phi[m]-\xi[m])/N_{\rm t}.$
@@
$\zeta_{3}\approx 2(1-\epsilon^{2})\epsilon^{2}\alpha_{k}(1-\alpha_{c})\phi[m].$ (33)
@@
$\zeta_{4}\approx 2(1-\epsilon^{2})\epsilon^{2}\alpha_{k}^{2}\phi[m].$ (34)
@@
$\mathrm{Var}\left\{X\right\}=(N_{\rm t}-K+1)\epsilon^{4}\alpha_{k}^{2}+\sum_{j\in\mathcal{K}}\alpha_{j}^{2}(1-\epsilon^{2})^{2}+\mu_{k},$ (36)
where $\mu_{k}$ is:
$\mu_{k}=2\alpha_{k}^{2}\epsilon^{2}(1-\epsilon^{2})(N_{\rm t}-K+1).$ (37)
It is worth noting that the term $\mu_{k}$ corresponds to the power of the term $X_{\rm{CC}}$. This completes the proof.